OpenStudy (anonymous):

Check all polynomials that are perfect square trinomials: 49x2 − 8x + 16 4a2 − 20a + 25 25b2 − 5b + 10 16x2 − 8x + 1

Parth (parthkohli):

Factor them and check if they are in the form \((\cdots)^2\)

OpenStudy (anonymous):

\[ 49 x^2-8 x+16=49 x^2-8 x+16\\ 4 a^2-20 a+25=(2 a-5)^2\\ 25 b^2-5 b+10=5 \left( 5 b^2-b+2\right)\\ 16 x^2-8 x+1=(4 x-1)^2 \] So second and third

OpenStudy (anonymous):

sorry second and fourth